what is the first number of a arithmetic sequence if the next two numbers are 3 and 17
Answer by Edwin McCravy(20056)

what is the first number of a arithmetic sequence if the next two numbers are 3 and 17

Let the first term be x:

The sequence goes

x, 3, 17

2nd term - 1st term = 3rd term - 2nd term
    3    -     x    =    17    -    3
          
Get the equation all together:

                3-x = 17-3
                3-x = 14
                 -x = 11
                  x = -11

The sequence is -11, 3, 17  

Answer: -11
             
Edwin